OpenStudy (anonymous):

Chris can be paid in one of two ways. Plan A is a salary of $460 per month, plus a commission of 8% of sales. Plan B is a salary of $688 per month, plus a commission of 5% of sales. For what amount of sales is Chris better off selecting plan A?

OpenStudy (anonymous):

Turn into equations. Call sales are S. Equation for: Plan A salary = ? Plab B salary = ?

OpenStudy (anonymous):

i dont understand the question

OpenStudy (anonymous):

i mean i cant find out how to get the equations out

OpenStudy (anonymous):

Simply put yourselves in the sales person's shoes. Your boss says I can pay you based on two different equations. Plan A and Plan B. Both depend on sales. The question is asking at what level of sales is Plan A better.

OpenStudy (anonymous):

Pay according to plan A = $460 + 8% of sales. = $460 + 0.08*S Pay according to Plan B = $688 + 5% of Sales = $688 + 0.05*S At what level of S will they be equal?

OpenStudy (anonymous):

sheg - you are of no help by throwing a number. goal is to educate so the person becomes good at solving problems. not taking answers.

OpenStudy (anonymous):

ok bro will remember this by the way answer is $ 7600

OpenStudy (anonymous):

do do i subtract 460 and 688 over and divide but the percentage but wouldnt that give me a - number?

OpenStudy (anonymous):

228 = 0.03*s s = 228/0.03 = $ 7600

OpenStudy (anonymous):

now im more confused...

OpenStudy (anonymous):

why did you find the difference of the 2 plans?

OpenStudy (anonymous):

The question is, at what level of S will plan A produce better pay. So, $460 + 0.08S must be greater than $688 + 0.05S. 0.08S - 0.05S > $688 - $460 => 0.03S > 228 => S > $288 * 100/3 => S > $7600 So, Plan A produces better when sales exceed $7,600. Given the two choices, sales people who are less confident of generating more than $7,600 sales will take plan B (pays more up front). Others will choose plan A.